OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y > -3x + 3 y > x + 2 (2, -5) (-2, 5) (2, 5) (-2, -5)

OpenStudy (amistre64):

your simplest bet is just to double check the given options

OpenStudy (anonymous):

funny i could have sworn we were just doing that http://openstudy.com/study?login#/updates/503b6e14e4b02c1631bd9bd8

OpenStudy (anonymous):

Is it C.

OpenStudy (anonymous):

it usually is C but lets check

OpenStudy (anonymous):

BRB

Parth (parthkohli):

lol "it usually is C"

OpenStudy (anonymous):

\(y > -3x + 3\) if \(x=2,y=5\) gives \[5>-3\times 2+3\] or \[5>-3\] that one works

OpenStudy (anonymous):

i should keep track, but i am going to say the vast majority of multiple guess turns out to be C one thing for sure, it is greater than \(\frac{1}{4}\) i would say it is close to \(\frac{5}{8}\) or even \(\frac{3}{4}\)